17. in organizing its incorporation in January, chupatch incurred the following costs: state incorporation fees $250, attorney fees: $2500 and promotional cost $3200. what journal entry would you make to record these expenses?

18. in their partnership agreement,Jane, Jill and Bridget have an income/loss distribution ratio of 5:3:2. how much would each of the three partners receive from an income of $120,000?

Jane will get $ :

Jill will get $ :

and Bridget will receive :
Answer #1


17: Journal Entry

incorporation fees $250
attorney fees $2500
promotional cost $3200
Cash $5950

18: Share of Income/Loss

Given ratio of 5:3:2 and income of $120,000

Jane will get : 120000*5/10 = $60000

Jill will get : 120000*3/10 = $36000

Bridget will get : 120000*2/10 = $24000